Two forces act a point The magnitude of the Two forces are F, 2 newton and the measure of the angle between them is 60, if their resultant equal 2√3 newton,then F =newton


1
Expert's answer
2021-11-10T10:21:06-0500

Given: Two forces as 2 N and F N, angle between them is "60\degree"and their resultant is "2\sqrt{3}" N.

We know that,

"R=\sqrt{R_1^2+R_2^2+2R_1R_2\cos\theta}"

"\Rightarrow 2\sqrt{3}=\sqrt{2^2+F^2+2\times2\times F\cos 60\degree}
\\\Rightarrow 2\sqrt{3}=\sqrt{4+F^2+2F} \ [\because \cos 60\degree=\frac 1 2]
\\\Rightarrow 12=4+F^2+2F
\\\Rightarrow F^2+2F-8=0
\\\Rightarrow (F-2)(F+4)=0
\\\Rightarrow F=2,-4
\\\text{As the magnitude of force can't be negative }
\\So, F=2N"